OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x + 7y = -21

OpenStudy (anonymous):

set x =0 solve for y then set y and solve for x

OpenStudy (anonymous):

y=3-x/7

OpenStudy (anonymous):

to find any "X" intercepts, set y=0 x+7(0)=-21 x=-21 is your "X" intercept To find any "Y" intercepts, set x=0 0+7y=-21 y=-21/7 y=-3 is your "Y" intercept Solution: Y intercept: {-3} X intercept: {-21}

OpenStudy (anonymous):

so wat could the answer be

OpenStudy (anonymous):

set x = 0 you get y = -3 so your answer for the y intercept is (0,-3) then set y = 0 and solve for x you get x =-21 so your x-int is (-21,0)

OpenStudy (anonymous):

understand awab?

OpenStudy (anonymous):

oooh thanksss

OpenStudy (anonymous):

no problem do you get how I did it?